OpenStudy (anonymous):

Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = -3 -15x - 5y = -10

OpenStudy (anonymous):

Umm use subsitution method here :D Using this method we can subsitute y

OpenStudy (anonymous):

2x+5y=-3 (equation 1) -15x-5y=-10 (equation 2) ======== -13x =-13 x=-13/-13 x=1 Now put value of x in equation 1 2x+5y=-3 2(1)+5y=-3 2+5y=-3 5y=-3-2 5y=-5 y=-5/5 y=-1